OpenStudy (anonymous):

what's the solution set for 2/1 > or equal to x

OpenStudy (anonymous):



OpenStudy (anonymous):

or -2.67 > or equal to x

OpenStudy (anonymous):

it starts at 2 and two is the greatest, so can you get it now?

OpenStudy (anonymous):

uhm no I dont get it would be better if you could be alittle more descriptive please...

OpenStudy (anonymous):

I really dont know how to write solution sets at all.

OpenStudy (anonymous):

okay the solution set of x<12 is {12 to -infinity} its all values for x that make the statement true. so apply that to yours

OpenStudy (anonymous):

i dont get it I quit...

OpenStudy (anonymous):

think of it this way then.... list down ALL the values of x that will this inequality true: \(\large 2\ge x \) i'll start you off.... let's try x=0... is \(\large 2 \ge 0 \) ??? yes, so x=0 is a part of the solution... x=1... is \(\large 2 \ge 1 \) ??? yes, so x=1 is a part of the solution... x=1.5... is \(\large 2 \ge 1.5 \) ??? yes, so x=1.5 is a part of the solution... x= 2.... is \(\large 2 \ge 2 \) ??? yes, so x=2 is a part of the solution... x=2.1 is \(\large 2 \ge 2.1 \) ??? NO, so x=2.1 is NOT a part of the solution... now keep doing this until you have ALL the x values that will make the statement \(\large 2 \ge x \) true.. keep in mind i only gave 5 numbers as an example... i don't have the time list down ALL of the x's that will make that statement true.... hmmm... there must be a shortcut way of writing 'em all down....